# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 4.5 Formulating the energy cascade in mass space
where the rate of energy cascade $\varepsilon_{u}$ can be directly related to the rate of mass cascade $\varepsilon_{m}$ (see Eq. (15)). Here, $\langle K_{pv}\rangle=\overline{K_{pv}}(0,t)$ is the mean specific kinetic energy for all halo particles. Similarly to Eq. (20), since $\varepsilon_{u}$ is scale independent, with continuous injection of virial kinetic energy $K_{pv}$ at a constant rate of $\varepsilon_{u}$ on the smallest mass scale, the rate of the energy cascade equals the rate of change of $\langle K_{pv}\rangle$ in all haloes. Therefore, we expect the total virial kinetic energy of all halo particles to be proportional to time $t$ or $\langle K_{pv}\rangle\propto t$. This is consistent with the cosmic energy evolution in Eq. (7).
5 Energy cascade in spherical haloes
5.1 Basic concepts
The mass and energy cascade in halo mass space describes the interactions between haloes of different masses. At this level, a statistically steady state is established to continuously maximize system entropy and release system energy while the N-body system evolves toward the limiting equilibrium. A key feature of that statistically steady state is the scale-independent rates of the mass and energy cascade ($\varepsilon_{m}$ and $\varepsilon_{u}$), as we demonstrate in Figs. 3 and 8.
Next, on the level of individual haloes, there exists a radial flow and radial velocity dispersion in dark matter, as shown by various simulations (Klypin et al., 2016; Wojtak et al., 2005). Similarly, mass and energy flow also exist along the halo radial direction, which is associated with the radial flow. In analogy to the cascade in the halo mass space, this section describes the mass and energy flow in haloes that is particularly important for the internal structures of haloes (or density profile) (Xu, 2023b), the universal scaling laws of dark matter haloes (Xu, 2023c), and the properties of dark matter particles (Xu, 2022b). The same theory can also be applied to the evolution and dynamics of the galaxy and associated scaling laws for the mass and size of bulges (Xu, 2024).
Figure 9: Schematic plot for the mass and energy flow along the radial direction in dark matter haloes. Both kinetic and potential energy increase with scale $r$. The mass accretion and self-gravity of haloes give rise to the inflow of an infinitesimal mass $dm_{r}$A" (green color). The outflow of an infinitesimal mass $dm_{r}$B" (red color) is due to the Hubble expansion. On a sufficiently small scale $r$, the inflow should balance the outflow such that the net mass flux vanishes on scale $r$. However, the energy flux can be finite. The outflow of "B" leads to a decrease in the total energy of "B", while the inflow of "A" leads to an increase in the total energy of "A". Kinetic energy flows from small to large scales ("inverse") at a rate of $\varepsilon$, accompanied by potential energy flowing from large to small scales ("direct") at a rate of $2/n_{e}\varepsilon$. The total energy flows from large to small scales at a rate of $(1+2/n_{e})\varepsilon$.
To understand the mass and energy flow in haloes, we start from the stable clustering hypothesis (SCH), a fundamental assumption for the nonlinear gravitational collapse at small scales. In principle, Hubble expansion leads to an outward radial flow and mass flux, while the halo mass accretion and self-gravity of haloes lead to an inward radial flow and mass flux. The SCH assumes that on sufficiently small scales, clusters of mass particles are bound and stable with a fixed mean physical separation between particles (Peebles, 1974; Davis and Peebles, 1977). There is no stream motion between particles in physical coordinates. In this sense, the peculiar motion cancels out the Hubble flow such that the inflow balances the outflow, and the net mass flux vanishes. However, the energy flux can be finite due to the non-uniform gravitational potential in haloes.
Figure 9 describes the mass and energy flow across a small scale $r$, where $v_{r}^{2}$ is the typical velocity dispersion on that scale. According to the stable clustering hypothesis, the structures below scale $r$ are bound and stable with constant mass and energy. During an infinitesimal time $dt_{r}$, the inflow of an infinitesimal mass $dm_{r}$ across scale $r$ due to the halo gravity (denoted by "1" in green color) balances the outflow due to the Hubble expansion (denoted by "2" in red color). The two infinitesimal masses can be far from each other without direct interactions to exchange energy or momentum. The mass flux into the scale $r$ should equal the mass flux out of the scale $r$, such that the net mass flux vanishes. Mathematically, the mass flux $\dot{m}_{r}$ and the infinitesimal time $dt_{r}$ read
$\begin{split}&\dot{m}_{r}\equiv\frac{dm_{r}}{dt_{r}}=\dot{m}_{r}^{out}=\dot{m}_{r}^{in}\quad\textrm{and}\quad dt_{r}=\frac{dr}{v_{r}}.\\ \end{split}$ (41)
The net mass flux vanishes, i.e., $\dot{m}_{r}^{net}=\dot{m}_{r}^{in}-\dot{m}_{r}^{out}=0$.
Next, let us consider the energy change on a scale of $r$. There are kinetic and potential energies associated with infinitesimal mass $dm_{r}$, both of which increase with $r$ and are related to each other by the virial theorem ($2KE=n_{e}PE$). Here, $n_{e}$ is an effective exponent and $n_{e}=-1$ for standard gravity and $n_{e}=-10/7$ from cosmic energy evolution (Eq. (7)). The potential and kinetic energies read
$\begin{split}KE=\frac{1}{2}{dm_{r}}{v_{r}^{2}}\quad\textrm{and}\quad PE=\frac{1}{n_{e}}{dm_{r}}{v_{r}^{2}}.\\ \end{split}$ (42)
Note that these energies are from the intra-halo interactions between particles in the same haloes. Only these energies are relevant to the energy flow in haloes. This is why we need to decompose the total particle energy into intra-halo and inter-halo contributions (Eq. (36)).
In Fig. 9, for an infinitesimal mass $dm_{r}$ (denoted as "A") at location $r+dr/2$ moves inward to $r-dr/2$ during time $dt_{r}$, the inflow (green arrow) due to the halo self-gravity leads to an increase in total energy (PE+KE) of infinitesimal mass $dm_{r}$,
$\begin{split}&\Delta{E}_{1}=-\frac{1}{2}\left(1+\frac{2}{n_{e}}\right)d{m}_{r}dv_{r}^{2}.\end{split}$ (43)
Now we can introduce a key parameter $\varepsilon^{*}$ to represent the rate of change in the specific energy of the infinitesimal mass A
$\begin{split}\varepsilon^{*}=\frac{\Delta{E}_{1}}{dm_{r}dt_{r}}=-\frac{1}{2}\left(1+\frac{2}{n_{e}}\right)\frac{dv_{r}^{2}}{dt_{r}}.\end{split}$ (44)
This is the key parameter describing the flow of total energy to small scales. Since the total energy is the sum of kinetic and potential energy, kinetic energy flows from small to large scales (inverse cascade with $\varepsilon_{KE}<0$), while the potential energy flows from large to small scales (direct cascade with $\varepsilon_{PE}>0$),
$\begin{split}\varepsilon_{r}=\varepsilon_{KE}=-\frac{1}{2}\frac{dv_{r}^{2}}{dt_{r}}=\frac{n_{e}}{2+n_{e}}\varepsilon^{*}\quad\textrm{and}\quad\varepsilon_{PE}=\frac{2}{2+n_{e}}\varepsilon^{*}.\end{split}$ (45)
In the same figure, for an infinitesimal mass $dm_{r}$ (denoted as "B") at $r-dr/2$ moving outward to $r+dr/2$, the outflow (red arrow) due to the Hubble expansion leads to a decrease in the total energy of the infinitesimal mass B
$\begin{split}&\Delta{E}_{2}=\frac{1}{2}\left(1+\frac{2}{n_{e}}\right)d{m}_{r}dv_{r}^{2}.\end{split}$ (46)
That decrease in the energy of mass $dm_{r}$ is "dissipated" due the the Hubble expansion, as we demonstrate in the cosmic energy evolution (Fig. 1). The rate of decrease in specific energy (the rate of energy "dissipation) of that $dm_{r}$ reads
$\begin{split}&\varepsilon_{a}=\frac{\Delta{E}_{2}}{dm_{r}dt_{r}}=\frac{1}{2}\left(1+\frac{2}{n_{e}}\right)\frac{dv_{r}^{2}}{dt_{r}}=-\varepsilon^{*},\end{split}$ (47)
which must balance the energy decrease on a cosmic scale (Eq. (8)). Since the energy decrease in mass B is due to the Hubble expansion (leaving the system as if "dissipated"), this leads to a continuous energy flow from large to small scales (the inflow of mass A). Clearly, the energy decrease ($\varepsilon_{a}$) of mass B is compensated by the energy flux ($\varepsilon^{*}$) due to the inflow of mass A such that the structures below scale $r$ maintain constant mass and energy.
Next, an important hypothesis we need to make is that the key parameter $\varepsilon_{r}$ for the flow of kinetic energy is independent of the scale $r$. While these haloes evolve toward limiting equilibrium, a statistically steady state is established. A key feature of this statistically steady state is a scale-independent rate for the energy flow in these haloes. Of course, this hypothesis must be tested and confirmed by simulations, as we demonstrate in Sections 5.3 (Fig. 11). At this time, we take this hypothesis to further illustrate the meaning of $\varepsilon_{r}$. Since $\varepsilon_{r}$ is independent of $r$, we can write $\varepsilon(m_{h},z)=-\varepsilon_{r}$ that can be a function of halo mass and redshift only. Solution of Eq. (45) gives the two-thirds law (with $dt_{r}$ from Eq. (41))
$v_{r}^{2}=\left(3\varepsilon r\right)^{2/3}.$ (48)
Integrating $dt_{r}=dr/v_{r}$ in Eq. (41), the typical time $t_{r}$ reads
$t_{r}=\int_{0}^{r}\frac{dr}{v_{r}}=\frac{3}{2}\frac{r^{2/3}}{(3\varepsilon)^{1/3}}=\frac{3}{2}\frac{r}{v_{r}}=\frac{v_{r}^{2}}{2\varepsilon}.$ (49)
Now, the rate of energy flow in haloes can be rewritten as
$\varepsilon=\frac{v_{r}^{2}}{2t_{r}}=\frac{v_{r}^{3}}{3r}.$ (50)
With changes in the specific energy from $v_{r}^{2}/2$ to zero during time $t_{r}$, the parameter $\varepsilon$ represents the rate of change in the specific energy of dark matter particles. Since the velocity dispersion $v_{r}^{2}$ increases with mass $m_{h}$, larger haloes tend to have a larger rate of energy flow.
Combined Eq. (48) with the virial theorem $v_{r}^{2}=\gamma_{r}Gm_{r}/r$ ($\gamma_{r}\approx 0.5$ is a numerical factor), we have
$m_{r}=\frac{1}{3\gamma_{r}}G^{-1}\varepsilon_{r}^{-1}v_{r}^{5}\quad\textrm{and}\quad\frac{dv_{r}^{2}}{dm_{r}}=\frac{2}{5}\frac{v_{r}^{2}}{m_{r}},$ (51)
where $m_{r}$ is the total mass enclosed in all scales below $r$. Finally, we can express $\varepsilon_{r}$ as (from Eqs. (48) and (51))
$\varepsilon=\frac{1}{2}\frac{dv_{r}^{2}}{dt_{r}}=\frac{1}{5}\frac{v_{r}^{2}}{m_{r}}\frac{dm_{r}}{dt_{r}}\quad\textrm{or}\quad\varepsilon m_{r}=\frac{1}{5}\frac{dm_{r}}{dt_{r}}v_{r}^{2}=\frac{1}{5}\dot{m}_{r}v_{r}^{2},$ (52)
where $\dot{m}_{r}=dm_{r}/dt_{r}$ is the mass flux at scale $r$. From Eq. (52), the product $\varepsilon m_{r}$ describes the energy flux associated with the mass flux $\dot{m}_{r}$, that is, the infinitesimal energy of ($dm_{r}v_{r}^{2}$) is transferred across the scale $r$ during infinitesimal time $dt_{r}$. The mass flux $\dot{m}_{r}$ and the energy flux $\varepsilon m_{r}$ are clearly related to the key parameter $\varepsilon$.
5.2 Potential fluctuations and random walk in haloes
In the previous section, we illustrate the mass and energy flow in haloes and identify a key parameter $\varepsilon$ to quantify the rate of flow. This section discusses the mass and energy flow on the particle level via random walk of particles in haloes. The random walk of haloes in halo mass space facilitates the mass cascade in mass space and leads to the halo mass function (Section 4). Similarly, the random walk of dark matter particles in haloes facilitates the energy cascade in haloes and leads to the halo density profile.
Figure 10: Schematic plot for the random walk of dark matter particles in haloes due to Hubble expansion. Particles conduct a random walk in a fluctuating gravitational potential field $\Phi=\Phi_{r}+E_{v}$, where $\Phi_{r}$ is the mean potential and $E_{r}$ is potential fluctuation. There is an energy flux associated with every elementary jump of a particle from small to large $r$ (position "1" to "2"), where the particle kinetic energy $v_{r}^{2}/2$ is cascaded to a larger scale $r$ after every migration. For a given particle, migration occurs with an average waiting time $\tau_{gr}(r,m_{h},z)$. Because particles have finite kinetic and potential energy, continuous migration facilitates an inverse kinetic energy cascade from small to large scales.
Figure 10 presents a schematic diagram for the random walk of particles in haloes with a fluctuating gravitational potential. Due to the discrete particle nature of haloes, the gravitational potential $\Phi$ in the haloes can never be smooth, as shown by N-body simulations. The total potential at scale $r$ can be decomposed into the mean and fluctuations. Here, $\Phi(r)=\Phi_{r}(r)+E_{r}(r)$, where $\Phi_{r}$ is the mean potential at scale $r$, and $E_{r}$ is the fluctuation of the potential. The mean potential, potential fluctuation, and density at scale $r$ read
$\begin{split}\Phi_{r}(r)=-\frac{Gm_{r}}{r},\quad E_{r}(r)=\frac{Gm_{p}}{d_{r}}\quad\text{and}\quad\rho_{r}(r)=\frac{m_{p}}{d_{r}^{3}}.\end{split}$ (53)
where $m_{r}$ is the mass enclosed by $r$, $d_{r}$ is the mean spacing between particles at scale $r$, and $m_{p}$ is the mass of a single particle.
Next, we focus on the average waiting time for the random walk of particles in a fluctuating potential, for example, from position "1" to "2" in Fig. 10. The local temperature $T_{r}$ and the probability $P_{r}$ for a successful attempt reads
$\begin{split}&k_{B}T_{r}=m_{p}v_{r}^{2}=m_{p}\gamma_{r}\frac{Gm_{r}}{r}\quad\textrm{and}\quad P_{r}=\exp\left(-\frac{m_{p}E_{r}}{k_{B}T_{r}}\right),\\ \end{split}$ (54)
where $v_{r}$ is the typical velocity at scale $r$ and is related to the local potential $\Phi_{r}$ by virial theorem. Hereafter, we adopt the subscript ’r’ to represent quantities on scale $r$, the subscript ’p’ for these quantities on the smallest scale in haloes that is determined by the particle mass and dependent on the resolution of the N-body simulations, and the subscript ’h’ for these quantities on the scale of halo size $r_{h}$.
Since the potential fluctuations are usually much smaller than the mean potential ($m_{p}E_{r}\ll k_{B}T_{r}$), the probability for a successful attempt and the waiting time $\tau_{gr}$ for particle random walk read
where $\tau_{gr}(r)$ is a position-dependent waiting time for the random walk of a given particle at location $r$. Here, $\Gamma_{r}(r)$ is the attempt frequency at $r$ (similarly to the attempt frequency in solid diffusion), i.e., the number of attempts per unit of time for a particle to jump to a neighboring position. The waiting time $\tau_{gr}$ can be related to the waiting time $\tau_{g}$ for the random walk of haloes in mass space (Eq. (18)), i.e. $\tau_{g}=\tau_{gr}(r=r_{h})$.
These two equations lead to
where $m_{h}^{*}$ and $r_{h}^{*}$ are the mass and size of haloes and $u_{h}^{*}=\Gamma_{r}d_{r}$ represents the speed of growth of haloes. Here, we use the fact that $m_{h}^{*}\propto t$, $r_{h}^{*}\propto t$, and $\tau_{g}^{*}\propto t^{0}$ (Eq. (24)). For a typical characteristic mass $m_{h}^{*}\approx 3\times 10^{13}M_{\odot}$, the size $r_{h}^{*}\approx 1$Mpc at $z=0$, and the age of the universe of 13.7 Billion years, the velocity $u_{h}^{*}$ is around 80km/s. Since the speed of halo growth $u_{h}^{*}$ remains as a constant while the density of characteristic haloes $\rho_{h}^{*}\propto a^{-3}$ decreases with cosmic time, we estimate that the $u_{h}^{*}$ is independent of halo density and should also be roughly independent of the scale $r$. The larger the mean particle distance $d_{r}$, the smaller the attempt frequency $\Gamma_{r}$. Therefore, the waiting time of the particles is inversely proportional to the local potential, that is, $\tau_{gr}\propto\Phi_{r}^{-1}$ from Eq. (55). Similarly, the waiting time of haloes in mass space is inversely proportional to the halo potential, i.e., $\tau_{g}\propto\Phi_{h}^{-1}$ in Eq. (18).
Next, consider a thin spherical shell of thickness $s_{r}$ located at $r$ that contains $N_{r}$ particles. We will identify the waiting time $\tau_{hr}(r)$ for a successful jump for any one particle of these $N_{r}$ particles in that spherical shell. In analogy to Eq. (9), we should have $\tau_{gr}(r)=N_{r}\tau_{hr}(r)$ and the waiting time
where $s_{r}$ is the thickness of the spherical shell or the jumping length of particle random walk. Similarly to Eq. (10) in mass space, the corresponding rate of mass flow due to the outward migration of particles should read (outflow in Fig. 9)
i.e. a mass of $m_{p}$ is transferred to larger scales during the time $\tau_{hr}$ via particle migration. Compared to Eq. (41), the particle mass $m_{p}$ is equivalent to the infinitesimal mass $dm_{r}$, while the waiting time $\tau_{hr}$ is equivalent to the infinitesimal time $dt_{r}$.
In analogy to Eqs. (49) and (50), we introduce a key parameter $\varepsilon$ (unit: m2/s3) to describe the rate of the energy flow along the halo radial direction, which takes the form of
$\begin{split}t_{r}=\frac{5}{2}dt_{r}\frac{m_{r}}{dm_{r}}=\frac{5}{2}\tau_{hr}\frac{m_{r}}{m_{p}}\quad\textrm{and}\quad\varepsilon=\frac{v_{r}^{2}}{2t_{r}}=\frac{v_{r}^{2}}{5\tau_{hr}}\frac{m_{p}}{m_{r}},\end{split}$ (60)
where $t_{r}$ is the characteristic time on scale $r$. Here, we use the scaling $t_{r}\propto v_{r}^{2}$ and $m_{r}\propto v_{r}^{5}$ (Eq. (51)). With a particle on scale $r$ migrating from position "1" to "2" (Fig. 10) during the waiting time $\tau_{hr}$, the specific energy of $m_{p}v_{r}^{2}/m_{r}$ is transferred across scale $r$. Similar to Eq. (52), the product $\varepsilon m_{r}$ in Eq. (60) represents the corresponding energy flux associated with the particle migration at scale $r$.
So far, we have only considered the outward mass flow as a result of the particle migration caused by the Hubble expansion, which leads to an increase in halo size $r_{h}$, a decreasing halo density with time, and energy "dissipation" (Eq. (46)). At the same time, the halo mass accretion (particles accreting onto the outskirts of haloes) and self-gravity lead to an increase in halo mass and introduce an inward radial flow as a result of the gravitational collapse (Fig. 10). At any radius $r$, the inflow is of the order $v_{r}$, the typical velocity at $r$ (Eq. (54)) such that the mass flow caused by the inflow reads
$\begin{split}\dot{m}_{r}=4\pi r^{2}\rho_{r}v_{r}.\end{split}$ (61)
For individual haloes, the competition between the migration-induced outward flow (Eq. (59)) and the accretion-induced inward flow (Eq. (61)) might lead to a non-vanishing net mass flux. However, a vanishing met mass flux is expected for the average radial flow for all haloes of the same mass, as required by the stable clustering hypothesis. In that case, we equate the out-flow flux in Eq. (59) and in-flow flux in Eq. (61), the thickness $s_{r}$ and waiting time $\tau_{hr}$ read
In this section, in analog to the halo migration in mass space, we introduce the particle migration in the fluctuating halo potential field and the associated mass and energy flux. The rate of energy flux $\varepsilon$ and the waiting time $\tau_{gr}$ and $\tau_{hr}$ are the central quantities for the energy cascade in the haloes that determine the internal structure, density profile, and many other properties of haloes. The energy cascade in haloes establishes a statistically steady state to continuously release the halo energy and maximize the halo entropy, where the rates of the energy cascade must be scale independent ($\varepsilon$ is independent of $r$). This leads to a scale-independent cascade between scales $r_{p}<r<r_{s}$, where the smallest scale $r_{p}$ depends on the particle mass or simulation resolution, while $r_{s}$ is the usual scale radius. The mathematical formulations are presented in Sections 5.3.
5.3 Formulating the energy cascade in haloes
This section quantifies the energy cascade in haloes from N-body simulations. The same theory can also be applied to the energy flow in galaxy bulges and the associated scaling laws for bulge mass, size, dynamics, and the evolution of the supermassive black holes (Xu, 2024).
In analogy to the formulation of the energy cascade in halo mass space (Section 4.5), we examine the energy cascade in all haloes of a given mass $m_{h}$ such that the averaged mass flux vanishes. We start the formulation by introducing the cumulative functions along the halo radial direction $r$. The cumulative mass function $\Lambda^{h}_{m}(m_{h},r,z)$ (similar to $\Lambda_{m}$ in Eq. (17)) represents the total mass above $r$ that is averaged for all haloes of the same mass $m_{h}$
$\Lambda^{h}_{m}(m_{h},r,z)=\int_{r}^{\infty}\rho_{r}\left(m_{h},r^{\prime},z)\right)4\pi r^{\prime 2}dr^{\prime}.$ (63)
where $\rho_{r}$ is the average mass density for all haloes of mass $m_{h}$.
Next, similarly to the formulation in Section 4.5, we also decompose the velocity of the halo particles $\boldsymbol{\mathrm{v}}_{p}$ into the halo mean velocity, $\boldsymbol{\mathrm{v}}_{h}=\langle\boldsymbol{\mathrm{v}}_{p}\rangle_{h}$, and the velocity fluctuation, $\boldsymbol{\mathrm{v}}_{p}^{{}^{\prime}}$, that is, $\boldsymbol{\mathrm{v}}_{p}=\boldsymbol{\mathrm{v}}_{h}+\boldsymbol{\mathrm{v}}_{p}^{\prime}$ (Eq. (36)). Here, $\boldsymbol{\mathrm{v}}_{h}$ represents the velocity of that halo, the average velocity of all particles in the same halo. Consequently, the total kinetic energy $K_{p}$ of a given halo particle can be divided into $K_{p}=K_{ph}+K_{pv}$. The virial kinetic energy, $K_{pv}={\boldsymbol{\mathrm{v}}_{p}^{\prime}}^{2}/2$, is the contribution from the velocity fluctuation due to the intra-halo interactions on small scales that are in the nonlinear regime. Only this part of the kinetic energy is relevant for the energy cascade in haloes. We introduce a cumulative function $\Lambda^{h}_{pv}$ for $K_{pv}$
As shown in Fig. 9, the inflow canceling out the outflow leads to a net vanishing mass flux. However, there exists a flux of specific kinetic energy $\varepsilon$ from small to large scales that leads to the increase in the particular energy on all scales greater than $r$, i.e., $\varepsilon=\Pi^{h}_{pv}$ (Fig. 9).
In a certain range of scales $r<r_{s}$ (inner haloes), the characteristic time on small scales is very small compared to the time on large scales. The small-scale motion does not feel the large-scale motion directly, except through the energy flux $\varepsilon$ across scales. When a statistical equilibrium is established, similarly to the energy cascade in the halo mass space, we expect the rate of energy flow $\varepsilon(m_{h},z)$ to be independent of the scale $r$, i.e.
This hypothesis can be tested by N-body simulations.
Figure 11: The variation of the rate of energy cascade $\Pi^{h}_{pv}(m_{h},r,z)$ (Eq. (66)) with radial scale $r$ at different redshifts $z$ for haloes of characteristic mass $m^{*}_{h}(z)$ and $m_{h}=0.044m^{*}_{h}(z)$. A scale-independent (independent of $r$) constant rate of $\varepsilon\equiv\Pi^{h}_{pv}(r<r_{s})$ can be identified for an inverse cascade of virial kinetic energy ${K_{pv}}$ from halo center to larger scales. There also exists a simultaneous direct cascade of potential energy from large to the smallest scale at a rate of $-7/5\varepsilon$ (Xu, 2022b). For haloes with characteristic mass $m^{*}_{h}$, the rate of energy cascade $\varepsilon(m_{h}^{*},z)\equiv\varepsilon_{u}\approx 10^{-7}m^{2}/s^{3}$.
where $\varepsilon_{u}\equiv\varepsilon(m_{h}^{*},z)$ is the rate of energy flow in haloes of characteristic mass $m_{h}^{*}$. The dimensionless parameter $\nu$ is defined as $\nu=({m_{h}}/{m_{h}^{*}})^{2/3}$ (Xu, 2023b). The rate of the energy cascade varies with the halo mass and time as $\varepsilon\propto m_{h}^{2/3}a^{-1}$, while the haloes evolve towards the limiting equilibrium. We only have a vanishing energy cascade or $\varepsilon=0$ in fully virialized haloes with $a\rightarrow\infty$ (usually small and old haloes), a limiting state that real haloes can never reach.
In the picture of energy flow in individual haloes, particle migration leads to an outward mass flow from small to large $r$. Halo mass accretion and self-gravity lead to an inward radial flow from large to small $r$. The two mass fluxes cancel such that the net mass flux vanishes. In addition, there is a continuous flow of kinetic energy from small to large scales with an energy flux $\varepsilon$ that is independent of the scale $r$ in a certain range of scales. A simultaneous flux of potential energy exists in the opposite direction. Again, according to the virial theorem ($02KE+PE=$ 0), the rate of a cascade of potential energy from large to small scales is approximately twice the rate of kinetic energy or $-2\varepsilon$. The more accurate calculation from cosmic energy evolution shows a rate of $-7/5\varepsilon$ for the potential energy with $n_{e}=-10/7$ in Fig. 9 (Eq. (7)).
5.4 Universal scaling laws and inner density of haloes
Scaling laws are well known to be associated with the energy cascade in turbulence (Richardson, 1922; Kolmogoroff, 1941). In this section, we focus on the scaling laws associated with the energy cascade in spherical haloes. The starting point is the expression of $\varepsilon$ (Eq. (50))
Since the rate of the energy cascade $\varepsilon$ is independent of $r$ in a certain range of scales, the 2/3 law can be easily obtained for kinetic energy $v_{r}^{2}\propto(-\varepsilon r)^{2/3}$. With virial theorem ($v_{r}^{2}\propto Gm_{r}/r$), the 5/3 scaling law can be recovered for the mass enclosed within $r$, i.e., $m_{r}\propto r^{5/3}$. We can write the scaling laws for mass $m_{r}$ (mass enclosed within $r$), density $\rho_{r}$ (mean density of the halo enclosed within $r$), velocity $v_{r}$, time $t_{r}$, and kinetic energy $v_{r}^{2}$, and mass flux $\dot{m}_{r}$ on scale $r$, all determined by three quantities $\varepsilon$, $G$ and scale $r$:
$\begin{split}&m_{r}=\alpha_{r}\varepsilon^{2/3}G^{-1}r^{5/3}\textrm{,}\quad\rho_{r}=\beta_{r}\varepsilon^{2/3}G^{-1}r^{-4/3},\\ &v_{r}^{2}=\beta^{-2/3}(\varepsilon r)^{2/3}\textrm{,}\quad t_{r}=0.5\beta^{-2/3}\varepsilon^{-1/3}r^{2/3},\\ &v_{r}^{2}=\beta^{-2/3}\alpha_{r}^{-2/5}(\varepsilon Gm_{r})^{2/5},\quad\dot{m}_{r}=4\pi\beta^{-1/3}\beta_{r}\varepsilon G^{-1}r\end{split}$ (70)
where $\alpha_{r}$ and $\beta_{r}$ are two numerical factors that can be determined by data fitting (Fig. 13). The evolution of these quantities is absorbed into $\varepsilon(m_{h},z)$. The relations between numerical factors are
$\begin{split}\beta=\frac{1}{3},\quad\gamma_{r}\alpha_{r}=\beta^{-2/3},\quad\textrm{and}\quad\beta_{r}=\frac{5}{4\pi}\alpha_{r}\beta\end{split}$ (71)
Other relevant quantities include the mean particle distance $d_{r}$, the attempt frequency $\Gamma_{r}$, the waiting time $\tau_{gr}$ and $\tau_{hr}$, and the shell thickness $s_{r}$. These quantities depend on the particle mass $m_{p}$ and the speed of growth for haloes of characteristic mass $m_{h}^{*}$, i.e., $u_{h}^{*}=\Gamma_{r}d_{r}\approx$80 km/s. From Eqs. (55) to (62),
$\begin{split}&d_{r}=\beta_{r}^{-\frac{1}{3}}m_{p}^{\frac{1}{3}}\varepsilon^{-\frac{2}{9}}G^{\frac{1}{3}}r^{\frac{4}{9}}\textrm{,}\quad\Gamma_{r}=\beta_{r}^{\frac{1}{3}}u_{h}^{*}m_{p}^{-\frac{1}{3}}\varepsilon^{\frac{2}{9}}G^{-\frac{1}{3}}r^{-\frac{4}{9}},\\ &\tau_{gr}=\frac{m_{p}}{\alpha_{r}\gamma_{r}}{u_{h}^{*}}^{-1}\varepsilon^{-\frac{2}{3}}Gr^{-\frac{2}{3}}\textrm{,}\quad\tau_{hr}=\frac{3}{5}\beta\gamma_{r}m_{p}\varepsilon^{-1}Gr^{-1},\\ &N_{r}=\frac{5{u_{h}^{*}}^{-1}}{3\alpha_{r}\beta\gamma_{r}^{2}}\varepsilon^{\frac{1}{3}}r^{\frac{1}{3}},\quad s_{r}=\frac{{u_{h}^{*}}^{-1}}{\alpha_{r}\beta^{1/3}}m_{p}\varepsilon^{-\frac{1}{3}}Gr^{-\frac{1}{3}}.\end{split}$ (72)
Since these quantities involve the particle mass $m_{p}$, information on these quantities may provide useful insights into the dark matter particle mass and properties.
[…]

### v4

@@ 3.4 The scale-to-scale energy cascade in mass space
where the rate of energy cascade $\varepsilon_{u}$ can be directly related to the rate of the mass cascade $\varepsilon_{m}$ (see Eq. (46)). Here, $\langle K_{pv}\rangle=\overline{K_{pv}}(0,t)$ is the mean specific kinetic energy for all halo particles. Similarly to Eq. (39), since $\varepsilon_{u}$ is scale independent, with continuous injection of virial kinetic energy $K_{pv}$ at a constant rate of $\varepsilon_{u}$ on the smallest mass scale, the rate of the energy cascade equals the rate of change of $\langle K_{pv}\rangle$ in all haloes. Therefore, we expect the total virial kinetic energy of all halo particles to be proportional to time $t$ or $\langle K_{pv}\rangle\propto t$. This is consistent with the evolution of cosmic energy in Eq. (122).
4 Energy cascade and halo density profile
In this section, we analyze the cascade of mass and energy operating within individual dark matter haloes. Because the dark component is collisionless, it does not support turbulent cascades driven by viscous dissipation and vorticity; similarly, a halo that has reached full virial equilibrium cannot sustain a net transfer of mass or energy across spatial scales. Accordingly, we use the term "cascade" to denote the scale‑to‑scale redistribution of mass and energy that occurs in nonequilibrium haloes as they relax toward that limiting equilibrium, rather than to imply a viscous, Kolmogorov‑type turbulent cascade. Analogously to the cascade and halo random walk in halo mass space, we characterize the intra‑halo mass and energy fluxes that play a central role in shaping halo internal structure (i.e., the density profile) (Xu:2023-Dark-matter-halo-mass-functions-and), underpinning the universal scaling relations of dark‑matter haloes (Xu:2023-Universal-scaling-laws-and-density-slope), and informing constraints on the properties of dark matter particles (Xu:2022-Postulating-dark-matter-partic). The same theory may also be applied to the evolution of the galaxy, yielding scalings for the mass and size of the bulge (Xu:2024-Cosmic-quenching-and-scaling-laws).
4.1 Particle random walk and distribution in haloes
Similarly to the halo mass function, the halo structure also depends on both initial conditions and gravitational dynamics. A more negative index or steeper spectrum usually leads to haloes with lower concentrations and less dense central cusps. Using Eq. (14) for asymptotic spectrum index $n_{a}$, a simple relation between the asymptotic density $r^{-\gamma_{a}}$ and the initial spectrum index $n_{i}$ can be obtained as
[…]
Here, we assume that the slope of the halo density is equal to the slope of the two-point correlation function $\xi(r)\propto r^{-(3+n_{a})}$. For the ratio parameter $\alpha_{x}=0$, this result reduces to the standard suggested result in Syer:1998-Dark-halo-mergers-and-the-formation.
Due to the inside-out nature of halo assembly, central regions collapse at high redshift from peaks in the primordial density field. Their structure is therefore largely set by the properties of the initial high-density peak and retains a stronger memory of the shape of the primordial power spectrum than do outer regions. By contrast, the outskirts are assembled later through mergers and smooth accretion; these processes drive violent relaxation, phase mixing, and tidal stripping, which progressively erase the imprint of the initial spectrum and render the dynamics near a characteristic scale $r_{s}$ predominantly universal. Consequently, the density profile near $r_{s}$ is expected to be approximately universal, governed primarily by gravitational dynamics, and only weakly sensitive to the initial spectrum.
To elucidate this behavior, consider an idealized initial condition in which the proto-halo is highly concentrated, with its mass localized at rr=0, representing the collapse of an extreme high-redshift peak. In this limit, the influence of the primordial spectrum is confined to a central boundary, whereas the subsequent evolution proceeds subject to self-gravity in an expanding background. This initial setting resembles the spherical collapse model (SCM). However, SCM enforces a uniform (top-hat) density and thus cannot capture the density variation. We instead adopt an effective mean-field kinetic description without considering gravity between individual particles. In contrast, dark matter particles move within a self-consistent gravitational potential field determined by the instantaneous particle distribution, permitting analytical halo density profiles.
The evolution of the particle distribution is modeled as a three-dimensional continuous-time random walk with a radius-dependent residence (waiting) time. In the gravity-dominated regime, this residence time is controlled by the local potential $\tau_{gr}(r)\propto\Phi_{r}(r)^{-1}\propto r^{-\gamma}$ (Eq. (75)), which admits an interpretation in terms of an effective local temperature: deeper potentials correspond to higher effective temperatures, larger particle kinetic energies, and shorter residence times. Around scale $r_{s}$, migration is dominated by gravity, yielding $\gamma=2/3$ (Eq. (58)). When the primordial spectrum dominates, the exponent $\gamma$ is expected to be related to the effective spectral index of the initial power spectrum (Eq. (52)). In a smooth dark-matter background, stochastic particle random walk expands the halo size and facilitates accretion. With smooth mass accretion, mass growth increases the total mass of the halo while leaving its internal structural profile essentially unchanged.
Next, we need to derive the particle distribution due to the random walk of particles. The 3D particle random walk can be described by a Langevin equation for the particle position $\boldsymbol{X}_{t}$,
[…]